please help me solve this quadratic question because i forgot how to do it
x^2 - 6 = 0

Answer by stanbon(75887)

x^2 - 6 = 0
---
Factor:
(x-sqrt(6))(x+sqrt(6)) = 0
----
x = sqrt(6) or x = -sqrt(6)
===============================
Cheers,
Stan H.